Opened 8 years ago

Closed 6 years ago

#437 closed flight processing (fixed)

GB12/07, flight day 088/2012, Ireland

Reported by: knpa Owned by: knpa
Priority: alpha 4 low Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: adbe, jaho Other processors:

Description (last modified by knpa)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/GB12_07-2012_088_Ireland

Data arrived from ARSF via network transfer on 28/03/2012

GB12/07 Scientific objective: The theme is to develop and verify techniques to assess and describe temporal soil moisture change though the cycle of the seasons. 4 flights are therefore planned, 2 in wet conditions and two in dry.

GB12/07 Priority: 7

PI: David Hughes

Sensors:

Camera (20/04/2012)
LiDAR (01/06/2012)
FW LiDAR

Change History (16)

comment:1 Changed 8 years ago by adbe

Navigation processed

comment:2 Changed 8 years ago by knpa

Added basestation data that PI made available via dropbox:

ASRF-3-12-GRGROUND-CONTROL-GAP

comment:3 Changed 8 years ago by adbe

Processed RCD, RCD ready for delivery check.

comment:4 Changed 8 years ago by knpa

  • Description modified (diff)

comment:5 Changed 8 years ago by jaho

RCD delivery check:

Everything looks good. Ready to deliver.

comment:6 Changed 8 years ago by adbe

  • Description modified (diff)

RCD delivered to David Hughes on 20/04/2012 on disk 122.

comment:7 Changed 7 years ago by adbe

Processed LiDAR, LiDAR ready for delivery check.

comment:8 Changed 7 years ago by adbe

  • Owner set to adbe
  • Status changed from new to accepted

comment:9 Changed 7 years ago by adbe

  • Cc adbe jaho added

comment:10 Changed 7 years ago by jaho

LiDAR delivery check:

All looks good. Ready to deliver.

comment:11 Changed 7 years ago by adbe

  • Description modified (diff)

LiDAR delivered to David Hughes on 01/06/2012 on disk 124.

comment:12 Changed 7 years ago by knpa

  • Description modified (diff)

comment:13 Changed 7 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from adbe to knpa
  • Status changed from accepted to assigned

Beginning archiving.

comment:14 Changed 7 years ago by knpa

Made available to NEODC.

comment:15 Changed 7 years ago by knpa

Archived by NEODC. Moved to archive disk.

comment:16 Changed 6 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed
Note: See TracTickets for help on using tickets.